OpenStudy (snitchseeker1496):

A proton enters a magnetic field of lux density 1.5 W b/m^(2) with a velocity if 2 x 10 ^(7) m/s at an angle of 30 degrees Celsius with the field. Compute the force on the proton.

OpenStudy (irishboy123):

you know this \({\mathbf {F}}=q{\mathbf {v}}\times {\mathbf {B}} \, \hat n\) ??? for the magnitude \( |\mathbf {F}| =q{ |\mathbf {v}} | \, |{\mathbf {B}}| \sin \theta \) \(q = 1.60217.....×10^{−19} C\) make sure you get \(\theta \) right way round. use the right hand rule if there is a drawing...